OpenStudy (anonymous):

INSTANT MEDAL which of the following is equivalent to 2s + 3 + s + 1 + 22 ? F. s + 2 = 22 G. s + 4 = 22 H. 3s + 2 = 22 I. 3s + 4 = 22

OpenStudy (anonymous):

2s + s = ? 3+1 +22=?

OpenStudy (anonymous):

s + 4 + 22!

OpenStudy (anonymous):

s + 4 = 22 i meant

OpenStudy (anonymous):

no.add the numbers that have letters with numbers that have letters with and numbers with no letters the other numbers with no letters

OpenStudy (anonymous):

so its 3s + 4 = 22 ! @Luis_Rivera